Please help me chose the correct answer.
Given sin(θ)=(3sqrt(17))/17, use the Pythagorean identity to find all possible values for tan(θ).
a) -sqrt(34)/17
b) sqrt(34)/17
c) -(2sqrt(34))/17
d) -(3sqrt(2))/4
e) (3sqrt(2))/4
f) (2sqrt(34))/17

Answer by Alan3354(69443)

Given sin(t)=3sqrt(17)/17, use the Pythagorean identity to find all possible values for tan(t).
-------------
cos(t) = sqrt(1 - sin^2) = sqrt(1 - 9/17) = 2sqrt(34)/17
--
tan = sin/cos
tan = 3sqrt(17)/(2sqrt(34)) = (3/2)/sqrt(2)
= 3sqrt(2)/4
The quadrant wasn't spec'd, so it's plus or minus.
a) -sqrt(34)/17
b) sqrt(34)/17
c) -(2sqrt(34))/17
d) -(3sqrt(2))/4
e) (3sqrt(2))/4
f) (2sqrt(34))/17